You have been managing a $5 million portfolio that has a beta of 1.20 and a required rate of return of 14%. The current risk-free rate is 4.75%. Assume that you receive another $500,000. If you invest the money in a stock with a beta of 1.15, what will be the required return on your $5.5 million portfolio? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

14=4.75+1.2*(market rate-4.75)

(14-4.75)=1.2*(market rate-4.75)

market rate=(14-4.75)/1.2+4.75

=12.4583333%(Approx)

Hence required return for $500,000=4.75+1.15*(12.4583333-4.75)

=13.6145833%(Approx)

Required return for portfolio=Respective return*Respective weight

=(5,000,000/5,500,000*14)+(500,000/5,500,000*13.6145833)

=13.96%(Approx).